# Survey on Modeling of Articulated Objects

### v1

@@ 4 Geometry Processing
  • part seg: whether the input shape of the object is pre-segmented into articulated parts with labels.
  • aligned: whether the object is aligned across different states when multiple articulated states are observed in the input; whether the objects are aligned in a canonical space when there are multiple objects in the input; for generative models, objects are assumed to be canonicalized by default.
  • # parts: whether the number of articulated parts is known.
  • partial/noisy: whether the input geometry can be partially observed (e.g. point cloud projected from the single view depth) or noisy (e.g. raw scans with sensory noise).
  • intermediate rep.: the intermediate representation used during the processing of the input shape or to represent the shape.
  • strategy: the methodology adopted for shape analysis, including handcrafted methods, supervised learning, and self-supervised learning. Handcrafted refers to the methods that rely on human-designed descriptors, heuristic rules, or mathematical algorithms to analyze the shape and it is not learning-based. Supervised refers to the data-driven methods that require ground truth labeled data for training, such as 2D masks, segmentation labels, displacement for each point, etc. Self-supervised refers to the data-driven methods that only rely on the input data for training, such as leveraging the geometric consistency between different states of the object.
  • supervision: the source of supervision used for the learning-based methods, including distance function (distance func.) typically used for metric learning, segmentation labels on the point cloud (labeled PC), segmentation mask on images (2D mask).
@@ 4.1.1 Articulated Part Perception
To ultimately analyze the part mobility of articulated objects, the recognition of the part segmentation from an input object shape is usually necessary as a first step. One exception to this is the early work from [MYY*10, ] and [HLV*17, ] that made the assumption that part segmentation is given as input to simplify the problem. ScrewNet [JLCN21] analyzes the part mobility from a point cloud sequence without knowing the part structure, but they directly predict the articulation parameters without explicitly assigning the points to different parts. Apart from these, other works for articulated part perception address the segmentation problem jointly with the part mobility analysis.
Perception from meshes. As the most widely used geometric representation in the industry, polygonal meshes are naturally chosen by the early works [XWY*09, MYY*10, SHL*14, HLV*17] as the initial input representation to work with. These meshes can be obtained from CAD models that are carefully designed by the artists with part-level structure and sharp edges. It implies that the input geometry in these works is assumed to be complete and noise-free, which simplifies the problem by observing the object in a geometrically perfect condition in 3D. [MYY*10, ] and [HLV*17, ] further simplify the problem of part mobility analysis by assuming that the partition of the object into articulated parts is accessible as input to the system. [SHL*14, ] pioneered addressing the segmentation problem for objects and articulated parts from an indoor scene as a mesh. Their segmentation relies on the difference between the object and part pose being observed in multiple states. They did not make any assumptions on the mesh properties, such as self-intersection or non-manifoldness. It loosens the requirement of the input geometry to be perfect, but it still assumes that the object is observed in a complete form with part structure available. It is expensive to obtain such a delicate mesh model from the real world, especially when we try to replicate a real-world object with a complex shape and part structure.
@@
Perception from point clouds. Point clouds of the objects can be obtained from 3D scans or depth sensors, which are more accessible and easier to obtain than the polygonal meshes. This accessibility makes the point clouds a more practical choice for the input representation as it can be directly obtained from the real world. It encourages more works later that consume point clouds as input for articulated part perception. Leveraging the ease of acquisition, a line of work [LWL*16, YLX*16, SCZ21, JLCN21] take point clouds that observe a sequence of articulated motion of the object as input. These point cloud sequences are assumed to be incomplete and noisy, which makes the robustness of the motion fitting an additional critical requirement for the system. Another implied assumption is that the object is aligned in the same coordinate system across different states, which is a prerequisite for the motion fitting process. However, observing a sequence of articulated motion of an object is unnatural in the real-world setting for passively articulated objects as they cannot move on their own. Consequently, the motion-capturing process typically demands extensive human intervention, making it costly and labor-intensive. Instead of observing motion sequences, [YHL*18, ] propose to take a pair of point clouds of different articulation states as input (illustrated in). Their work also relaxes several assumptions made by the previous works that the object can be misaligned in two states and the input only needs to be from objects that are functionally related but not necessarily from the same instance. More recent works [WZS*19, YHY*19, LWY*20, LZH*23, DLS*23] tackle the problem from a single state of the object at inference time to make it more accessible on the input. [AFA*22, ], [LXX*22, ], and GAPartNet [GXZ*23] further extend the input to colored point clouds, which can leverage more visual information than the raw point clouds. One note for the work from [LSH*23, ] is that they conduct the articulated part analysis from an over-segmented point cloud, which is a different problem setting from the other works in this category.
Figure 4: Method proposed by [YHL*18, ] to analyze the articulated part structure from a pair of point clouds from functionally related objects in arbitrary articulation states. Figure reproduced from original paper[YHL*18]
Perception from monocular RGB images. Even more accessible than point clouds, monocular RGB(-D) images have also served as input for articulated part perception recently. [QJR*22, ] proposed to extract 3D planes to represent articulated parts from RGB videos recording the human-object interaction. OPD [JMSC22] first introduces the task of detecting openable parts of objects from a single-view image. OPDMulti [SJSC23] extends the scenario to multiple objects in the scene as a follow-up. The articulation analysis in 3D from a single-view RGB image is a challenging but useful task, especially for the robotics system equipped with RGB cameras.
@@ 4.1.2 Articulated Object Creation
Articulated object reconstruction. The goal of articulated object reconstruction is to recover the 3D geometry of the articulated objects with either articulation parameters estimated from input observations or the ability to animate the objects to different articulation states. Since triangle meshes are the most widely used geometric representation for 3D objects, it is naturally chosen as the output representation for the works in this category. [PG08, ] presents the earliest work that reconstructs the surface from a sequence of point clouds, where they assume the part segmentation and skeleton of the rigid pieces are given in the first frame. Once all the other frames are aligned with the first one by tracing the skeleton, the mesh surface of each part is accumulated and recovered using point cloud meshing algorithms [KBH06, ACTD07]. A-SDF [MQK*21] is the first work that reconstructs articulated objects from signed distance fields (SDFs) of a single snapshot in a way that the objects can be deformed to any other articulation states. [WCM*22, ] share the same goal with A-SDF but take multi-view RGB images as input for the reconstruction. CARTO [HIZ*23] further extends the input to a single stereo RGB image. One limitation of these three works is that they both reconstruct the articulated object as a whole into a single surface. The lack of part-level structure in the output geometry makes it difficult to actually interact with the object in a meaningful way. CLA-NeRF [TLYS22], Ditto [JHZ22], and PARIS [LMS23] address this limitation by reconstructing the articulated object in the part level. CLA-NeRF and PARIS both take multi-view RGB images as input, while Ditto consumes point clouds. Ditto and PARIS share the configuration that the object is observed from two articulated states and the two states are assumed to be pre-aligned. However, research in this category typically assumes a predetermined number of articulated parts, concentrating on modeling single moving parts or relying on the object’s category to infer its structure. Such assumptions highlight a notable gap in the literature: the challenge of reconstructing objects with an arbitrary number of articulated parts in varying structures remains insufficiently addressed.
Articulated object generation. This category of work aims to synthesize the 3D geometry of articulated objects with articulation parameters compatible with each part. This line of work falls into the category of generative models, which typically learn the distribution of the input data and generate new samples from it. NAP [LDS*23] is a pioneering work that generates a full description of the articulated object, including part geometry, articulation graph, and joint parameters from random noises unconditionally. CAGE [LTMS23] extends the task in a conditional setting, where the object category and a part connectivity graph as given as the user-driven input. The output of these works targets the mesh representation composited in varied structures. The input for these works is not explicitly specified, but it is assumed that the input is aligned and noise-free, which is a common assumption for generation tasks.
@@ 4.2 Methodology
Figure 5: A selection of the intermediate representation used for articulated part perception. Figures reproduced from original papers [LWY*20, LWL*16, MST*21].
@@ 4.2.1 Intermediate Representation
Articulated part perception. For the work focusing on articulated part perception, the intermediate representation plays an important role in the methodology to facilitate the analysis of the part mobility by providing a more informative and expressive representation of the input geometry. Below we discuss each representation used for this category.
@@
  • Surface patch. The methods taking mesh as input commonly leverage the surface patches as the proxy for analysis. The local patches can be detected from slippable analysis as kinematic surfaces [GG04], which are used to associate with rigid motions for each part [XWY*09]. The surface patches can also be useful for revealing the self-similarity and symmetry on the mechanical part to characterize the motion configuration [MYY*10]. The patches at the intersection area between components can be leveraged to determine the supporting-supported relationship [SHL*14], which are informative for the part decomposition and structure analysis. Similarly, the interaction pattern described at the connecting region between surface patches is leveraged in [HLV*17, ] to measure the similarity between snapshots of the object for motion classification.
@@
  • 3D trajectory. The 3D trajectory is a spatial-temporal representation that records the position of the points in the 3D space over time. For works capturing motion with a sequence of point clouds as input, extracting the 3D trajectory is a common choice for the intermediate representation [PG08, YLX*16, SCZ21]. As the points are initially unordered and untracked, establishing the correspondence between the points across different frames is a critical step for motion fitting. Once the trajectory is robustly estimated, the rigid pieces in the object can be grouped based on the transformational consistency of the points.
  • NAOCS, NPCS. Articulation-aware Normalized Coordinate Space Hierarchy (ANCSH) consists of two coordinate spaces: Normalized Articulated Object Coordinate Space (NAOCS) and Normalized Part Coordinate Space (NPCS). It is a two-level hierarchical representation that was first proposed by [LWY*20, ]. It is designed manually for articulated objects to align the instances and parts to a canonical space. NAOCS provides a canonical space on the object level that binds with a specific object category, while NPCS provides a canonical reference for each articulated part from a specific object instance. Knowing the object category as input, mapping the input shape to the NAOCS will produce part segmentation as it is defined by the structure template for each object category. The object category is not only a semantic label but also a structural label that defines the spatial arrangement of the parts. For example, a cabinet stacked with two drawers is classified separately from a cabinet with three drawers. Mapping the segmented parts to the NPCS can be useful for estimating the articulated state or pose for each part. Inspired by this idea, [LXX*22, ] and GAPartNet [GXZ*23] propose to use NPCS as the intermediate representation for the articulated part analysis.
  • Point cloud feature (PC feat.). Point cloud latent features are commonly used as the intermediate representation in the methods that take point clouds as input. The learning-based methods typically extract useful features in 3D from the input point cloud and use them to assign the part label and to regress the articulation parameters. PointNet [QSMG16], PointNet++ [QYSG17], and Sparse U-Net [GEvdM18] are the common feature extractors backboned in the existing works.

### v3

@@ 4 Geometry Modeling
  • # parts: as an assumption, whether the number of articulated parts is known in the input object.
  • partial/noisy: whether the input geometry can be partially observed (e.g. point cloud projected from the single view depth or even without depth) or noisy (e.g. raw scans with sensory noise).
  • aligned: as an assumption, whether the object is pre-aligned across different states when multiple articulated states are observed in the input; whether the objects are aligned in a canonical space when there are multiple objects in the input; whether the objects are assumed to be canonicalized for generative models.
  • part seg.: whether the methods rely on precomputed part segmentation as the input assumption.
  • intermediate rep.: the intermediate representation used during the processing of the input shape or to represent the shape. Please refer to Section 4.2.1 and Figure 7 for more explanation.
  • strategy: the methodology adopted for shape analysis, including handcrafted methods, supervised learning (SL), semi- and weakly supervised learning (semi-weakly SL), transfer learning (TL), and self-supervised learning (SSL). Handcrafted refers to non-learning-based methods that rely on human-designed descriptors, heuristic rules, or mathematical algorithms to analyze the shape. SL refers to the data-driven methods that require ground truth labeled data for training, such as segmentation labels, displacement for each point, etc. SSL refers to the data-driven methods that only rely on the input data for training, such as leveraging the geometric consistency between different states of the object. Semi-weakly SL refers to a hybrid learning paradigm that combines elements of semi-supervised learning and weakly supervised learning to leverage both limited strongly labeled data, abundant weakly labeled data, and unlabeled data for model training. TL refers to the data-driven methods that adapt pre-trained foundational models for a specific target task or use a pre-trained model as a feature extractor to adapt to a related task.
  • supervision: the source of supervision used for the learning-based methods, such as segmentation labels on the point cloud (labeled PC), segmentation mask on images (2D mask), signed distance field (SDF) or occupancy field (occupancy) of the object, multi-view RGB(-D) images paired with camera parameters (MV RGBs or MV RGB-Ds), and unposed RGB images (RGBs).
  • arti. part seg.: the shape of the object is segmented into articulated parts as the output.
@@ 4.1.1 Articulated Part Perception
To analyze the part mobility of articulated objects, segmenting articulated parts from an input shape is typically a necessary step. Exceptions are the early work by mitra2010illustrating and hu2017learning which simplified the problem by assuming articulated part segmentation was provided as input, and ScrewNet jain2021screwnet bypassing explicit segmentation by directly analyzing mobility from a point cloud sequence. Beyond these, other articulated part perception methods jointly address the articulated part segmentation and mobility analysis from unstructured input shapes, with the exception of liu2023semi which first leverages fine-grained semantic segmentation to group articulated parts for motion analysis and later uses them as weak labels to enhance perception performance. In the following, we categorize existing works based on the geometric representation they take as input and discuss the assumptions made about the input data.
Figure 5: Method proposed by sharf2014mobility for object/part segmentation from a mesh input by relying on the objects/parts being observed in multiple states or poses. Figure reproduced from original paper sharf2014mobility.
Perception from meshes. As the most widely used geometric representation in the industry, polygonal meshes are naturally chosen by early works xu2009joint; mitra2010illustrating; sharf2014mobility; hu2017learning as the initial input representation to work with. These meshes can be obtained from CAD models that are carefully designed by the artists with part-level structure and sharp edges. It implies that the input geometry in these works is assumed to be complete and noise-free, which simplifies the problem by observing the object in a geometrically perfect condition in 3D. xu2009joint takes a complex mesh model as input and segments the model into components for joint analysis used for articulate or deform the model. The proposed method is demonstrated on a range of models, including connected multi-component models, disconnected multi-component models, and single-component models. mitra2010illustrating focus on the motion illustration for mechanical assemblies represented as a set of polygonal meshes that has been partitioned into individual parts. They assume that each part is modeled as a 2-manifold surface with no self-intersections beyond a small tolerance. sharf2014mobility pioneered addressing the mobility analysis for indoor scenes represented as unorganized boundary meshes. They did not make any assumptions on the mesh properties, such as self-intersection or non-manifoldness. However, segmenting objects and their parts from a scene requires observing state variations in the object and its components, as illustrated in Figure 5. hu2017learning learns part mobility model from a single snapshot of the object represented as polygonal meshes. It assumes that the part segmentation is given as input and the method applies on one pair of reference-moving parts at a time. A recent work by qiu2025anymesh proposes to convert a static mesh model into its articulated counterparts in a open-vocabulary manner. It does not assume any pre-computed part segementation and category information given as input. This approach leverages the recent advances in image segmentation models to assist the 3D segmentation and use VLMs for the part mobility analysis.
The main goal of these mesh-based works is to analyze the mobility on artist-created models or manufacturing designs, which can be potentially useful in automating the animation process, illustrating the motion of mechanical assemblies, or creating interactive simulations. Some challenges are commonly shared among these works, such as the complexity in processing varying mesh structures with irregular topology and the difficulty in converting the mesh into a more compact representation for learning-based methods. The assumption of having clean and complete geometry in the input also limits the applicability of these methods to real-world object reconstructions, which are often incomplete and noisy.
@@
Perception from point clouds. Point clouds of the objects can be obtained from 3D scans or depth sensors, which are more accessible and easier to obtain than the polygonal meshes. This accessibility makes the point clouds a more practical choice for the input representation as it can be directly obtained from the real world. It encourages more works later that consume point clouds as input for articulated part perception (examples shown in LABEL:fig:pcinput). Leveraging this ease of acquisition, a line of work li2016mobility; yuan2016space; shi2021self; shi2021self; jain2021screwnet; weng2021captra; nie2023sfa; liu2023building takes point clouds that observe a temporal sequence of articulated motion of the object as input. MultiBodySync huang2021multibodysync also processes multiple scans but does not assume a temporal sequence. These point cloud sequences are often incomplete and noisy, which makes the robustness of the motion fitting an additional critical requirement for the system. Another implied assumption is that the object is aligned in the same coordinate system across different states, which is a prerequisite for the motion fitting process. By registering the point positions over time, the points exhibiting consistent motion patterns are grouped as the articulated parts. The motion patterns can be further analyzed to estimate the articulation parameters for each part. However, observing a sequence showing object motion can sometimes be unnatural for objects that cannot move on their own. Consequently, the motion-capturing process typically demands extensive human intervention, making it costly and labor-intensive. Another challenge in the motion capturing process is how to filter out the irrelevant components in the observed sequence, such as the human hand or robotic arm.
Instead of observing motion sequences, yi2018deep and Ditto jiang2022ditto take a pair of point clouds of different articulation states as input. yi2018deep relaxes several assumptions made by the previous works that the object can be misaligned in two states and the input only needs to be from objects that are functionally related but not necessarily from the same instance. Ditto assumes the point cloud pair is pre-aligned from the same object and also assumes only one part is moving in the observation. But it produces the mesh surface with each part segmented as the output to build digital twins of the object. More recent works wang2019shape2motion; yan2019rpm; li2020category; liu2023self; deng2023banana tackle the problem from a single state of the object at inference time to make the input more accessible. abdul2022learning, liu2022toward, and GAPartNet geng2023gapartnet further extend the input to colored point clouds, which can leverage more visual information than the raw point clouds. In tackling the challenge of lacking training data with ground-truth labels, liu2023semi proposes to leverage data with fine-grained semantic segmentation from PartNet mo2019partnet as weak labels to augment the training data for the part mobility analysis.
Perceiving part mobility from point clouds facilitates the analysis of real-world objects with reduced domain transfer effort compared to mesh-based methods. Point clouds are easier to obtain from depth sensors, making them a practical choice for many applications. As a discrete approximation of object surfaces, they are flexible, lightweight, and easier to process with learning-based methods. However, their lack of connectivity information and fine surface details makes them less informative than meshes, which can capture intricate geometric features more effectively. This limitation poses a challenge for learning-based methods, as they must extract meaningful features from sparse and unstructured data to accurately understand part structures.
Perception from monocular RGB images. Even more readily available than point clouds, RGB(-D) images are another common input for articulated part perception. qian2022understanding proposed to extract 3D planes to represent articulated parts from RGB videos recording the human-object interaction. OPD jiang2022opd first introduces the task of detecting openable parts of objects from a single-view image. OPDMulti sun2023opdmulti extends the scenario to multiple objects in the scene as a follow-up. wang2025active further improves the accuracy of the articulated part segmentation from real images by leveraging active learning to easily obtain the ground truth labels and propose a new dataset for the task. Compared to the aforementioned methods that rely solely on 3D geometry, RGB-based approaches eliminate the need for extensive 3D capturing or reconstruction, making them highly beneficial for real-time applications. These methods leverage visual cues in images, providing richer appearance details and contextual information about the object. However, they also face unique challenges, including the difficulty of inferring 3D structures from 2D images, handling occlusions and scene clutter, and compensating for limited or missing 3D information in the input.
@@ 4.1.2 Articulated Object Creation
In recent years, significant advances in neural representation techniques mildenhall2021nerf; wang2021neus; kerbl20233d, along with large vision models oquab2023dinov2; kirillov2023sam; wang2024dust3r and generative models goodfellow2014gan; ho2020ddpm; rombach2022high; blattmann2023stable, have opened new possibilities for 3D modeling for articulated objects. These approaches enable more flexible and detailed representations of articulated objects, which have traditionally posed challenges due to their complex geometries and dynamic characteristics. This emerging direction in articulated object creation leverages the strengths of both neural geometry representations and generative models, facilitating innovative methods for reconstructing or generating articulated objects across various scenarios.
Articulated object reconstruction. The goal of articulated object reconstruction is to recover the 3D geometry and motion parameters of the objects from various observations. On the geometry side, the challenges not only involve decomposing the part into articulated components in the perceived format, but also producing the reconstructed surface that can be animated into different articulation states. The input modalities can be diverse, including point clouds, multi-view RGB or RGB-D images, and RGB videos. Certain assumptions are made about the input data based on the target object types and the sparsity of observations. These include factors such as the number of articulated parts, the object’s alignment across different states, and the completeness of the input data. The output representation is typically the mesh surface, which is widely used for animation and simulation. A line of work focuses on reconstructing the whole object as a single deformable surface either over time or given different states mu2021asdf; wei2022self; heppert2023carto; song2024reacto; swaminathan2024leia; kerr2024robot. Another line of work reconstructs each part as separate surfaces interconnected by joints to enable articulation tseng2022cla; jiang2022ditto; liu2023paris; weng2024neural; mandi2024real2code.
Figure 6: Representative methods for articulated object generation. Figure reproduced from original papers lei2023nap; liu2024cage; li2024dragapart; gao2024meshart; liu2024singapo; le2024articulate.
pekelny2008articulated pioneered surface reconstruction from sequential point clouds, assuming known part segmentation and skeleton in the first frame. Building on this idea of observing objects in motion, WatchItMove noguchi2022watch extends the input to multi-view RGB videos and approximates each moving part as an ellipsoid. RSRD kerr2024robot models objects in 4D space by taking multi-view RGB images of their static state and a casual RGB video demonstrating human interaction. Similarly, REACTO song2024reacto proposes to reconstruct 4D objects from a single causally captured monocular RGB video of the object in motion. While these methods accommodate arbitrary kinematic structures by registering different object states, they require dense motion observations. A-SDF mu2021asdf was the first to reconstruct articulated objects using signed distance fields from a single snapshot, allowing deformation into different articulation states. wei2022self share the same goal with A-SDF but take multi-view RGB images as input for the reconstruction, while CARTO heppert2023carto further extends the input to a single stereo RGB image. These works ease the data acquisition process by requiring only a single snapshot of the object, but they assume the object has only one movable part. LEIA swaminathan2024leia reconstructs the interpolated state of the object between the two given input states from multi-view images without knowing the number of moving parts. However, one limitation of the above-mentioned works is that they all reconstruct the articulated object as a single unified surface or as an implicit deformation field. The lack of part-level structure in the output geometry makes it difficult to interact with the object in a physically realistic manner.
Recent works address this limitation by reconstructing the articulated object in the part level. SfA nie2023sfa designs a pipeline to reconstruct real-world objects by iteratively interacting with the moving parts and capture the object in the point cloud sequence. To make the reconstruction more efficient, CLA-NeRF tseng2022cla proposes to reconstruct the object from a few multi-view RGB images of an unseen 3D object instance within the known category. It assumes that the object category is known and all the instances in the category share the same articulated structure. To avoid the reliance on object category, Ditto jiang2022ditto reconstructs from a pair of point clouds, while PARIS liu2023paris take multi-view RGB images observing the object in two states as input. weng2024neural extends the setting of PARIS to multi-view RGB-D images and is able to reconstruct multiple articulated parts in the output. These works can reconstruct the object across arbitrary categories, but they all assume the number of articulated parts is known at inference time. In a more flexible setting, Real2Code mandi2024real2code proposes to reconstruct articulated parts from multiple unposed RGB images of the object in an opening state. This work focuses on certain categories of objects that mostly have cuboid-like parts but with no assumption on the number of articulated parts. Capturing the opening state of the object is also helpful in modeling the interior geometry of the object, which is usually occluded in the closed state. However, this method relies on the object being observed in a state that maximally exposes each part, which may be constrained by its natural configuration or the observation angle.
Despite the rapid progress in articulated object reconstruction, existing methods still face several challenges. Early approaches primarily focused on surface reconstruction from sequential point clouds or dense motion observations but lacked explicit part-level structure, making realistic interaction difficult. Recent works have addressed this by reconstructing objects at the part level, but they often rely on strong priors such as known object categories, predefined articulation structures, or a fixed number of moving parts. This highlights a critical gap in the field: the need for methods that can reconstruct arbitrary articulated objects with unknown part structure from sparse and unconstrained observations.
Articulated object generation. This line of work aims to synthesize the articulated objects in both geometric and kinematic structures from random noise or user-specified constraints. The representative work is illustrated in Figure 6. This is a newly emerging direction of research that leverages the recent advances in neural representation learning and generative modeling in creating high-fidelity and diverse 3D assets. These generative models learn the underlying distribution of articulated objects, capturing articulated part geometry and articulation configurations to generate novel samples. NAP lei2023nap is a pioneering work that unconditionally generates a full description of an articulated object, including part geometry, articulation graph, and joint parameters. MeshArt gao2024meshart tackles the same task but designs a different geometry representation by directly generating the mesh vertices from the latent code. CAGE liu2024cage, SINGAPO liu2024singapo, and ArtFormer su2024artformer introduce different controllable generation setting: CAGE incorporates high-level structural information (e.g., articulation graph) as user constraints, SINGAPO enables single-image-driven generation, and ArtFormer focuses on text-driven generation. Instead of learning the distribution from scratch, ARTICULATE-ANYTHING le2024articulate proposes a general framework that automates the articulation of objects from different input modalities, including text, images, and videos. Programs are used to arrange the part geometries and predict articulation with a critic mechanism designed to ensure the plausibility of the generated object. The output of these works is a mesh representation that is either synthesized from the intermediate representation or retrieved from a database using the object context. A key challenge in this direction is representing articulated objects in a compact and expressive way to capture the joint distribution of part geometry and articulation effectively. Another challenge is enabling precise control over the generation process, especially for complex, highly structured objects.
On a different front, DragAPart li2024dragapart introduced a method to generate images of articulated objects in different states from a single input image, driven by user-specified drag actions. While the output remains in 2D space, its articulation is 3D-aware, allowing users to interactively explore the object’s motion in a 2D setting. This highlights an under-explored yet promising direction: simulating 3D-aware articulation dynamics within a 2D framework.
@@ 4.2.1 Intermediate Representation
Articulated part perception. In the task of articulated part perception, various intermediate representations are employed to support geometric analysis, either through neural networks or handcrafted methods. These intermediate representations serve as critical abstractions that simplify the complex structures of articulated objects, making them more tractable for further processing. Depending on the type of input data, these representations can take different forms, including surface patches for capturing local geometric details, 3D trajectories for tracking motion over time, point cloud features for representing spatial relationships, or image features for extracting visual cues from 2D data. Each representation is chosen based on the specific demands of the task, and it plays a crucial role in bridging the gap between raw perception data and high-level geometric reasoning.
@@
  • Surface patch. The early methods taking mesh as input commonly leverage the surface patches as the proxy for analysis. The local patches can be detected from slippable analysis as kinematic surfaces gelfand2004shape, which are used to associate with rigid motions for each part xu2009joint. The surface patches can also be useful for revealing the self-similarity and symmetry on the mechanical part to characterize the motion configuration mitra2010illustrating. The patches at the intersection area between components can be leveraged to determine the supporting-supported relationship sharf2014mobility, which are informative for the part decomposition and structure analysis. Similarly, the interaction pattern described at the connecting region between surface patches is leveraged in hu2017learning to measure the similarity between snapshots of the object for motion classification.
@@
  • 3D trajectory. The 3D trajectory is a spatial-temporal representation that records the position of the points in the 3D space over time. For early works capturing motion with a sequence of point clouds as input pekelny2008articulated; yuan2016space, explicitly extracting the 3D trajectory is a common choice as the intermediate representation to optimize the correspondences between points along the timeline. Once the trajectory is robustly estimated, the rigid pieces in the object can be grouped based on the transformational consistency of the points.
  • Point cloud feature (PC feat.). Point cloud latent features are commonly used as the intermediate representation in the methods that take point clouds as input. The learning-based methods typically extract spatial point-wise features from the input point cloud and supervise them for the specific objectives, such as part labeling, point registration, or articulation parameter estimation. PointNet qi2016pointnet, PointNet++ qi2017pointnet++, and Sparse U-Net graham2018submanifold are the common feature extractors used as a backbone method in existing works. For the works that takes point cloud sequence or pair as input yi2018deep; shi2021self; huang2021multibodysync; weng2021captra; liu2023building; lin2024autourdf, the point cloud features are used to capture point correspondence for temporal consistency across different states. For other works that take single state point cloud as input wang2019shape2motion; yan2019rpm; li2020category; liu2023self; deng2023banana; abdul2022learning; liu2022toward; geng2023gapartnet; liu2023semi, the point cloud features are used to encode the spatial relationships between points for articulated part labeling, while additional heads are usually added to predict the motion parameters or part pose along with the part segmentation.
  • Convex. The convex is used to approximate the input geometry by representing the part as a convex hull. liu2023few proposes to learn deformation networks by choosing convexes as intermediate representation to deform each part mesh to generate new objects by few-shot learning.